OpenStudy (anonymous):

Find the area bounded between the curve and the x-axis on the given interval. f(x) = sec(x) on [0, π/4]

OpenStudy (amistre64):

integrate sec(x) from 0 to pi/4

OpenStudy (amistre64):

sec(x) was actually integrated before the integration process was even concieved

OpenStudy (anonymous):

how can you tell if it was already integrated?

OpenStudy (amistre64):

some map maker was wanting to do something or another, and he worked out the formula needed to integrate it

OpenStudy (amistre64):

\[\int sec=\int sec\frac{sec+tan}{sec+tan}=\int\frac{sec^2+sec~tan}{tan~+~sec}\] now that the top is the derivative of the bottom; this is just the derivative of a log

OpenStudy (amistre64):

\[\int \frac{1}{u}du\]

OpenStudy (anonymous):

when you use the interval 0, pi/4 is it the smaller value minus the big value or vice versa?

OpenStudy (amistre64):

top minus bottom: big and small have no bearing on it \[\int_{a}^{b}f(x)~dx=F(b)-F(a)\]

OpenStudy (amistre64):

the interval from a to b ...

OpenStudy (anonymous):

i'm still not getting the answer. if it is 1/u du, then shouldn't it be ln(sec(pi/4))

OpenStudy (amistre64):

\[ln(\frac{tan(45)+sec(45)}{tan(0)+sec(0)})\] \[ln(\frac{1+\sqrt2}{0+1})\] \[ln(1+\sqrt2)\] is what i get

OpenStudy (amistre64):

with any luck i did it right :)

OpenStudy (amistre64):

there isnt some continuity issues in that interval is there?

OpenStudy (anonymous):

how'd you get the ln you got: ln (tan45 + sec45/ ....)

OpenStudy (amistre64):

well; i showed how to modify sec(x) to get a function that is of the from 1/u du u = tan+sec \[\int_{a}^{b}\frac1udu=ln(u(b))-ln(u(a))\] a property of logs is that subtraction is division \[ln(u(b))-ln(u(a))=ln(\frac{u(b)}{u(a)})\] \[ln(\frac{tan(b)+sec(b)}{tan(a)+sec(a)})\]